OpenStudy (anonymous):

integrate sin^2x / sqrt x dx

OpenStudy (anonymous):

Is it \[\int \frac{sin^2x}{\sqrt{x}}dx\]

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

this is a tough one..im trying integration by parts but i dont know

OpenStudy (dumbcow):

its not possible http://www.numberempire.com/integralcalculator.php